# Variable aggregation for nonlinear optimization problems

### v2

@@ 4.2 Moving bed reactor optimal operation
This reactor is modeled with nonlinear, one-dimensional differential-algebraic equations discretized along the length domain of the reactor, as described by Ostace et al. ostace2018 and Okoli et al. okoli2020 and implemented in the IDAES model library lee2021idaes . In this test problem, we penalize the deviation from a set-point of 95% oxygen carrier conversion and an inlet solid flow rate of 591 kg/s.
@@ 4.3 Pipeline network dynamic optimization
The gas pipeline network optimization problem (PIPE) is a constrained NLP consisting of a system of DAEs modeled in IDAES lee2021idaes . The goal is to minimize the gas transmission cost while meeting the customer demands and satisfying the mass and momentum balance equations for gas transport NAIK20231847 . IDAES provides a modular approach to model the gas pipeline flowsheet by leveraging existing unit models and material properties. However, this approach also leads to additional redundant variables and constraints that link the individual unit models and property packages to build the overall flowsheet.
@@ 4.4 AC optimal power flow
Alternating current optimal power flow (ACOPF) describes the task of selecting power generator dispatch levels and bus voltages to minimize cost while satisfying the AC power flow equations over a network cain2012 . We use as a test problem a 4,917-bus case from the PGLib-OPF test set pglib . The test problem data is parsed and converted into a Pyomo model by Egret knueven2019 .
@@ 5 Computational results
Runtime results are those obtained by solving a particular instance of a model with a particular solver (here, IPOPT), such as solve times and numbers of iterations. To limit the impact of difficult instances, which may have extremely long solve times, and to avoid questions about how to penalize instances that don’t converge, we compare detailed timing statistics only for nominal instances of each test problem with each aggregation method applied. Runtime results are obtained by timing evaluation callbacks in the CyIpopt interface to IPOPT cyipopt .
@@ 5.1 Structural results
We first compare the effect of variable aggregation on model structure. We define model structure as solver-agnostic properties of the optimization model that do not depend on particular variable values, such as numbers of variables and constraints. A summary of model structures after aggregating with each method is given in Table 3.
@@
Figure 5: Percent of variables eliminated by each method for each model. Methods are in the same order as presented in Table 3
@@ 5.2 Runtime results
Variable aggregation can also have a significant effect on an optimization problem’s solve time. We note that the direction of impact of variable aggregation on solve time is not immediately obvious. Aggregating variables leads to a smaller, denser optimization problem with potentially more inequality constraints (if many bounded variables are eliminated) and potentially more nonlinear constraints (if many variables are replaced with nonlinear expressions). The dominant computational costs in an interior point method are function and derivative evaluations and KKT matrix factorization. These are less expensive for a smaller problem, but KKT matrix factorization is more expensive for a denser problem, and Hessian evaluation is more expensive for a denser problem with more nonlinear constraints. For these reasons, it is important to study the breakdown of runtime to understand which factors are contributing to speedup or slow-down caused by variable aggregation. Finally, an interior point method may take significantly different numbers of iterations with full-space and aggregated optimization problems. While it is difficult to predict which formulation will converge in fewer iterations, it is important to observe trends that occur for a specific problem or problem class.
@@ 5.3 Convergence reliability
Convergence results of a parameter sweep over relative volatility and the feed mole fraction of heavy component for the distillation optimal control problem are shown, for each method, in Figure 9. Results of a parameter sweep varying inlet solid temperature and inlet solid flow rate in the moving bed model are shown in Figure 10 and results of a sweep varying gas temperature and supply pressure in the pipeline dynamic optimization problem are shown in Figure 11. A summary showing the percent of problems solved by each method, for each model and in total, is shown in Table 6.
@@
% conv $t_{solve}^{avg}$ conv $t_{solve}^{avg}$ conv $t_{solve}^{avg}$
@@ 6 Conclusion
Different variable aggregation strategies are compared across four test problems in terms of structure, solve time and convergence reliability. The aggregation strategies are grouped into two main categories:
@@
Different aggregation strategies aggregate significantly different number of variables across all test problems. The approximate maximum aggregation strategies, specifically the linear matching-based heuristic, eliminate a significantly higher percentage $(70-90\%)$ of variables compared to the structure-preserving strategies which eliminate $<60\%$ variables. However, it is important to note that the number of non-zeros per constraint stays approximately the same when structure-preserving aggregation strategies are used, while the approximate maximum aggregation strategies lead to an increase in the total number of non-zeros per constraint and also the increase the nonlinear non-zeros making the problem more nonlinear.
@@
Based on our results, we believe that approximate-maximum and incidence-preserving aggregation methods should be implemented as pre-solve options in nonlinear modeling environments and solvers. However, further explanation and confirmation of our results is necessary to determine whether these aggregation methods are sufficiently general to be applied by default. We believe that the following efforts should be undertaken to explain and confirm the trends we observe:
@@
    Theoretical analysis of interactions among solvers’ globalization methods and different aggregation strategies.
@@ Appendix A Block triangularization
2:$G_{d}=\text{\tt project}(G,\mathcal{M})$triangleright The nodes of $G_{d}$ are $A$
3:$\mathcal{C}=\text{\tt strongly\_connected\_comps}(G_{d})$triangleright $\mathcal{C}$ partitions $A$
4:$D=\text{\tt compress}(G_{d},\mathcal{C})$triangleright The nodes of $D$ are $\mathcal{C}$
5:$\mathcal{S}=\text{\tt topological\_sort}(D)$
6:$\mathcal{T}=\{\}$
7:for $C\text{~{}\bf in }\mathcal{S}$ do $\triangleright$ $C$ is a subset of $A$
8:     $T=\{\}$triangleright $T$ is a subset of $\mathcal{M}$
9:     for a in $\text{~{}\bf in }C$ do

### v3

@@ 4.2 Moving bed reactor optimal operation
This reactor is modeled with nonlinear, one-dimensional differential-algebraic equations discretized along the length domain of the reactor, as described by Ostace et al. ostace2018 and Okoli et al. okoli2020 and implemented in the IDAES model library lee2021idaes. The model consists of multilinear and rational nonlinearities along with terms with fractional powers. In this test problem, we penalize the deviation from a set-point of 95% oxygen carrier conversion and an inlet solid flow rate of 591 kg/s.
@@ 4.3 Pipeline network dynamic optimization
The gas pipeline network optimization problem (PIPE) is a constrained NLP consisting of a system of DAEs modeled in IDAES lee2021idaes. The goal is to minimize the gas transmission cost while meeting the customer demands and satisfying the mass and momentum balance equations for gas transport NAIK20231847. The gas network model consists of nonlinear terms with ratios and fractional powers to compute power consumption in the compressors IDAES provides a modular approach to model the gas pipeline flowsheet by leveraging existing unit models and material properties. However, this approach also leads to additional redundant variables and constraints that link the individual unit models and property packages to build the overall flowsheet.
@@ 4.4 AC optimal power flow
Alternating current optimal power flow (ACOPF) describes the task of selecting power generator dispatch levels and bus voltages to minimize cost while satisfying the AC power flow equations over a network cain2012. The model consists of sinesine and cosinecosine nonlinearities arising from the power flow equations. We use as a test problem a 4,917-bus case from the PGLib-OPF test set pglib. The test problem data is parsed and converted into a Pyomo model by Egret knueven2019.
@@ 5 Computational results
Runtime results are those obtained by solving a particular instance of a model with a particular solver (here, IPOPT). These include solve time and number of iterations. To limit the impact of difficult instances, which may have extremely long solve times, and to avoid questions about how to penalize instances that don’t converge, we compare detailed timing statistics only for nominal instances of each test problem with each aggregation method applied. Runtime results are obtained by timing evaluation callbacks in the CyIpopt interface to IPOPT cyipopt.
@@
Table 3: Summary of method names and algorithm numbers
Method name Shorthand name Algorithm Example acceptable constraint
Linear degree-1 LD1 3 $y=1$
Equal-coefficient degree-2 ECD2 6 $y=x$
Linear degree-2 LD2 5 $y=2x$
Degree-2 D2 4 $y=x^{2}$
Greedy GR 1 $y=f(w,x,\dots)$
Linear-matching LM 2 $y=f(w,x,\dots)$
@@ 5.1 Structural results
We first compare the effect of variable aggregation on model structure. We define model structure as solver-agnostic properties of the optimization model that do not depend on particular variable values, such as numbers of variables and constraints. A summary of model structures after aggregating with each method is given in Table 4. For reference, the shorthand name and algorithm number of each method can be found in Table 3.
@@
Figure 5: Percent of variables eliminated by each method for each model. Methods are in the same order as presented in Table 4.
@@ 5.2 Runtime results
Variable aggregation can also have a significant effect on an optimization problem’s solve time. We note that the direction of impact of variable aggregation on solve time is not immediately obvious. Aggregating variables leads to a smaller, denser optimization problem with potentially more inequality constraints (if many bounded variables are eliminated) and potentially more nonlinear constraints (if many variables are replaced with nonlinear expressions). The dominant computational costs in an interior point method are function and derivative evaluations and KKT matrix factorization. These are less expensive for a smaller problem, but KKT matrix factorization is more expensive for a denser problem, and Hessian evaluation is more expensive for a denser problem with more nonlinear constraints. For these reasons, it is important to study the breakdown of runtime to understand which factors are contributing to speedup or slow-down caused by variable aggregation. Finally, an interior point method may take a significantly different number of iterations with full-space and aggregated optimization problems. While it is difficult to predict which formulation will converge in fewer iterations, it is important to observe trends that occur for a specific problem or problem class.
@@ 5.3 Convergence reliability
Convergence results of a parameter sweep over relative volatility and the feed mole fraction of heavy component for the distillation optimal control problem are shown, for each method, in Figure 9. Results of a parameter sweep varying inlet solid temperature and inlet solid flow rate in the moving bed model are shown in Figure 10 and results of a sweep varying gas temperature and supply pressure in the pipeline dynamic optimization problem are shown in Figure 11. A summary showing the percent of problems solved by each method, for each model and in total, is shown in Table 7. The average solve time ($t^{avg}_{solve}$) in Table 7 is computed over the successfully converged instances of the parameter sweep.
@@
% conv $t_{solve}^{avg}$ % conv $t_{solve}^{avg}$ % conv $t_{solve}^{avg}$
@@ 6 Conclusion
A novel linear-matching variable aggregation strategy is introduced and compared with existing strategies across four nonlinear optimization problems. The strategies strategies are compared across four test problems in terms of structure, solve time and convergence reliability. The aggregation strategies are grouped into two main categories:
@@
Different aggregation strategies aggregate significantly different number of variables across all test problems. The approximate maximum aggregation strategies, specifically the linear matching-based heuristic, eliminate a significantly higher percentage $(70-90\%)$ of variables compared to the structure-preserving strategies which eliminate $60\%$ variables. However, it is important to note that the number of non-zeros per constraint stays approximately the same when structure-preserving aggregation strategies are used, while the approximate maximum aggregation strategies lead to an increase in the total number of non-zeros per constraint and also the increase the nonlinear non-zeros making the problem more nonlinear. We note that the success of the linear-matching strategy at eliminating large numbers of variables can also be leveraged to identify large nonsingular submatrices for Schur complement decompositions in linear systems, which is a potential avenue for future work.
@@
Based on our results, we believe that approximate-maximum and incidence-preserving aggregation methods should be implemented as pre-solve options in nonlinear modeling environments and solvers with an incidence-preserving aggregation method as the default option. Additionally, we believe that the following efforts should be undertaken to explain and confirm the trends we observe:
@@
    Theoretical analysis of interactions among solvers’ global convergence methods (e.g., line search and trust region methods) and different aggregation strategies.
@@ Appendix A Block triangularization
[…]
7:for $C\text{~\bf in }\mathcal{S}$ do $\triangleright$ $C$ is a subset of AA
8:  $\{\} \triangleright TT$ is a subset of $\mathcal{M}$
9:  for a $\text{~\bf in }C$ do